PROBLEM 4

The capital structure of the Valley Products Company is as shown below, on April 30, 2019.

Long term Debt (8%)​​ $ 128,000,000

Common Stockholders equity (8 million shares) 192,000,000

​​​ = $ 320,000,000

The company believes this capital structure to be optimal and therefore desires to maintain it. That is, DO NOT CHANGE IT. The company estimates next years net income, NII, will be $48 million. Its payout ratio is expected to remain at 25 percent.

New bonds will have a 10 percent coupon rate and will be sold at par. However, investment bankers agree that if Valley sells more than $30,000,000 in bonds, the cost of debt for the entire issue will rise to 13 percent. The required rate of return on common stock, rs, is 15.0%, but due to flotation cost, the firm must earn 18.0% on its new common stock, r. The firm's marginal tax rate is 40 percent.

4a) The firm wishes to maintain its current capital structure. If the capital budget for Valley Products is $67.5 million, will the firm use the cheaper 10 percent or the more expensive 13 percent debt in the capital budget? Support your answer.



4b. Remember, the firm wishes to maintain its current capital structure. If the capital budget for Valley Products is $67.5 million, will the firm use retained earnings or new common stock as the equity financing component in the capital budget? Support your answer.





4c. What is the cost of capital for Valley Products if its capital budget equals $67.5 million? Remember that the firm wishes tomaintain its current capital structure. The firm's marginal tax rate is 40%.
Answer #1
$ mn Proportion
Debt 128 40.00%
Equity 192 60.00%
Total 320 100.00%

Part 4a

Internal accruals = Net income x (1 - payout ratio) = 48 x (1 - 25%) = $ 36 million

Capital budget = $ 67.5 million

Incremental debt = 40% = 40% x 67.5 =  27.00 million

Equity required = 67.5 x 60% = $ 40.5 million

Since incremental debt of $ 27 million is less than the cut off debt of $ 30 million, hence the firm will use the cheaper 10 percent debt in the capital budget.

Part (b)

Incremental equity = $ 40.5 million > internal accruals of $ 36 million, hence the firm will use

  • Entire retained earnings of $ 36 million + new common stock = 40.5 - 36 = $ 3.5 million as the equity financing component in the capital budget

Part (c)

Cost of capital has been calculated as 12.32% as shown below:

Funded by $ mn Proportion Cost Tax rate Post tax cost Component Cost
A B C T D = C x (1 - T) C x D
Debt       27.00 40.00% 13% 40% 7.800% 3.1200%
Internal accruals (Retained earnings) as equity       36.00 53.33% 15% 15% 8.0000%
External equity          4.50 6.67% 18% 18% 1.2000%
Total       67.50 100.00% 12.3200%